OpenStudy (anonymous):

What is negative 3 5/7 as an improper fraction? A. -21/7 B. -7/26 C. -15/7 D. -26/7 @iGreen @TexasBoy123 @PaulaLovesSchool13 @ttb123456789 please help

OpenStudy (anonymous):

will help

OpenStudy (igreen):

Multiply the denominator to the whole number, then add it to the numerator. \(3 \dfrac{5}{7} \rightarrow \dfrac{7\cdot 3 + 5}{7}\)

OpenStudy (igreen):

Can you multiply and add 7 * 3 + 5?

OpenStudy (anonymous):

i dont unterstand problom i'm not an english speasker

OpenStudy (anonymous):

yes its 21+5=26 duh

OpenStudy (anonymous):

oh its D. thanks to the both of you

OpenStudy (igreen):

Yep, so that gives you \(\dfrac{26}{7}\).